OpenStudy (anonymous):

What are the solutions to the equation x2+4=148 Ill give you a medal :/

OpenStudy (jadeishere):

You want to isolate the variable so you have to move 4 over to the other side of the equality sign. x^2 + 4 = 148 -4 -4 x^2 = 144 x = the square root of 144 x = 12

OpenStudy (anonymous):

Thanks !

OpenStudy (jadeishere):

Yep :) Whatever you do to one side of the = sign, you do to the other, until you get one variable isolated